# RAGDiffusion: Faithful Cloth Generation via External Knowledge Assimilation

### v1

@@ Abstract
Standard clothing asset generation involves creating forward-facing flat-lay garment images displayed on a clear background by extracting clothing information from diverse real-world contexts, which presents significant challenges due to highly standardized sampling distributions and precise structural requirements in the generated images. Existing models have limited spatial perception and often exhibit structural hallucinations in this high-specification generative task. To address this issue, we propose a novel Retrieval-Augmented Generation (RAG) framework, termed RAGDiffusion, to enhance structure determinacy and mitigate hallucinations by assimilating external knowledge from LLM and databases. RAGDiffusion consists of two core processes: (1) Retrieval-based structure aggregation, which employs contrastive learning and a Structure Locally Linear Embedding (SLLE) to derive global structure and spatial landmarks, providing both soft and hard guidance to counteract structural ambiguities; and (2) Omni-level faithful garment generation, which introduces a three-level alignment that ensures fidelity in structural, pattern, and decoding components within the diffusing. Extensive experiments on challenging real-world datasets demonstrate that RAGDiffusion synthesizes structurally and detail-faithful clothing assets with significant performance improvements, representing a pioneering effort in high-specification faithful generation with RAG to confront intrinsic hallucinations and enhance fidelity.
@@ 1 Introduction
While the high quality of images generated by diffusion models [27, 57] has significantly lowered the barriers for individuals to engage in convenient content creation across various tasks, the generation of 2D standard clothing assets from in-the-wild data has been relatively underexplored. Standard clothing asset generation refers to generating forward-facing flat-lay garment images [78] on a clear background by recovering clothing information from arbitrary real-world scenes (e.g., garments hung on hangers, worn by models, or casually laid on chairs in Fig. 1). Standard garment serves as an important intermediary variable connecting multiple downstream applications, such as garment design, product display, and virtual fitting [24, 81, 80].
Standard clothing generation is a challenging high-specification generative task primarily due to the highly constrained manifold of standard clothing distributions [78] and the high sensitivity of users, which demands a strong fidelity to the generated results. In practice, we have observed that the quality of existing general generative models [52, 69] fails to reach a level that is realistically usable, especially concerning the structure of garments. Actually, these structurally unreliable outcomes, referred to as structure hallucinations [4], are commonly seen in the outputs of large generative models [48, 44], manifesting as distorted facial expressions, palms with 6 fingers, and nonsensical text [4], etc. Current generative models often prioritize compatibility with multi-modal or multi-conditional inputs [34, 49], sampling acceleration [57, 38], and zero-shot editing capabilities [16, 25]. A limited number of works [47, 20, 1] have attempted to mitigate the phenomenon of structural hallucinations through extensive training with a large number of trainable parameters. However, the unsustainable training costs and data requirements render this approach impractical [7]. In this study, we take a step towards addressing the prevalent structure hallucinations in the generation process, aiming for higher specification image generation in the future.
@@
Inspired by the recent advancements in retrieval-augmented NLP [9, 51], we propose RAGDiffusion, a more general and efficient solution, by introducing Retrieval-Augmented Generation (RAG) [51] to enhance structural determinacy during the generation process. Our insight is to eliminate the inherent uncertainty of generative processes by assimilating structural representations and knowledge from external models and databases. This approach avoids the traditional method of heavily compressing visual concepts into a large model through costly training, thereby alleviating the burden on generative models. RAGDiffusion, as a RAG paradigm for addressing the issue of structural illusions, includes two processes: information aggregation based on retrieval, and conditional generation with omni-level fidelity. (1) Retrieval-based Structure Aggregation: Contrastive learning [23, 15] is introduced to train a dual tower network to extract multi-modal structure embeddings from images of two branches as well as attributes derived from a large language model (LLM) [5, 3]. Considering potential encoding bias in real-world data, we propose a structure retrieval named Structure Locally Linear Embedding (SLLE) to project the predicted structure embedding towards a standard manifold [54, 70] as well as offer a silhouette landmark. The remapped latent structure embedding and the landmarks provide both soft and hard structure guidance respectively to eliminate structural hallucinations. (2) Omni-level Faithful Garment Generation: To enhance the faithfulness of the standard garment, RAGDiffusion involves three-level alignment into the generation process of the diffusion model to provide feedback on structure, pattern and decoding faithfulness for garment components. Specifically, we align spatial structure with Embedding Prompt Adapter and Landmark Guider, ensure the generated pattern detail matches the conditioning by ReferenceNet [28], and reduce distortion during reconstruction from VAE [32] by Parameter Gradual Encoding Adaptation (PGEA) to adapt the SDXL [48] backbone to a more powerful VAE.
To the best of our knowledge, RAGDiffusion stands as a pioneering multi-modal RAG method to solve intrinsic hallucination and unfaithfulness during image synthesis. Our innovative insight of incorporating external prior knowledge with retrieval to mitigate uncertainty has the potential to catalyze advancements across a broader spectrum of faithful generation applications as showcased in training-free demos in  Sec. 4.5. We summarize the contributions of this paper as follows:
@@
  • We propose RAGDiffusion, a novel RAG framework for standardized clothing generation with representation learning and SLLE to offer structure determinacy and eliminate hallucination.
@@
  • We employ a three-level alignment in RAGDiffusion, addressing structure, pattern, and decoding aspects, ensuring faithfulness of clothing on this high-specification task.
@@
  • Comprehensive experiments on challenging in-the-wild sets demonstrate RAGDiffusion is capable of synthesizing both structure and detail faithful clothing assets, outperforming current methods by a substantial margin. Furthermore, the ablation study has validated the effectiveness of different parts of RAGDiffusion.
@@ 2 Related works
Retrieval-augmented generative models. Retrieval-augmented strategies leverage external databases to enhance the capabilities of generative models across a variety of tasks. For instance, the RETRO [8] modifies an LLM to effectively utilize external databases, achieving impressive performance. In the realm of image synthesis, retrieval has been employed in both GANs [59, 12] and diffusion models [7, 56], accommodating 3D generation [55], video generation [76] and artistic styles [53]. The essence of these works lies in retrieving similar images to serve as mimetic references for generating specified content, particularly in cases where the generative models are insufficiently trained. In contrast, our RAGDiffusion aims to leverage retrieval to aggregate structural information, thereby mitigating intrinsic hallucinations and boosting the performance of large base generative models on high-standard tasks.
Controllable text-to-image diffusion models. To attain conditional control in text-to-image diffusion models, ControlNet [75], T2I-Adapter [43] and IP-Adapter [69] incorporate additional trainable modules to fuse conditions on feature maps. Additionally, recent investigations have utilized a variety of prompt engineering techniques [33, 67, 79] and implemented cross-attention constraints [14, 64, 30, 82] to facilitate more controllable generative processes. Furthermore, some studies further investigate multi-condition or multi-modal generation within a single diffusion model [29, 49, 83, 34]. However, these approaches rely on the associative capabilities of stable diffusion, which do not guarantee fidelity in terms of structure and detail in the generated results, particularly when dealing with the highly standardized sampling of standard garments.
@@
Image restoration. Recent studies  [60, 39, 71] have made pioneering strides in developing universal models for natural image restoration. These investigations encompass tasks such as weather effect removal, low-light image enhancement, denoising, and deblurring. One of the most popular applications is image inpainting [65, 48, 11, 26, 41] which utilizes text prompts to guide the generation of content within editing regions. Most similar to our task, document image restoration [73, 62, 68] addresses issues such as dewarping [40, 72, 17, 21], deshadowing [36, 74, 35], and illumination correction [73, 62, 18]. However, unlike 2D document restoration, standard garment generation is inherently a 3D semantic content restoration task and may be influenced by factors such as occlusions, accessories, and human body parts, presenting greater challenges in semantic localization. Consequently, existing image restoration techniques fall short of achieving faithful cloth generation.
@@ 3 Method
An overview of the RAGDiffusion is presented in Fig. 2. The backbone of RAGDiffusion employs the SDXL [52], with the preliminary detailed in Appendix. Given an in-the-wild clothing image $x_{itw}\in\mathbb{R}^{H\times W\times 3}$, RAGDiffusion is aimed to generate an authentic standard flat lay in-shop garment image $x_{std}$. A dual tower StructureNet extracts latent structure embeddings by contrastive learning as detailed in Sec. 3.1, and SLLE retrieves and fuses structure embeddings and landmarks as in Sec. 3.2. Three-level alignment generation involves Embedding Prompt Adapter, ReferenceNet, and PGEA for structure, pattern, and decoding faithfulness described in  Sec. 3.3.
@@ 3.1 Dual-tower embeddings extraction
Contrastive learning. Given a batch of $N$ (in-the-wild clothing $x_{itw}$, standard clothing $x_{std}$) pairs, StructureNet $g_{\theta}$ is trained to predict which of the $N\times N$ possible ($x_{itw}$, $x_{std}$) pairings across a batch actually occurred according to structure similarity. To do this, StructureNet learns a multi-modal embedding space by jointly training a dual tower encoder tailored for $x_{itw}$ and $x_{std}$ to maximize the cosine similarity of the embeddings of the $N$ real pairs (marked with superscript $+$) in the batch while minimizing the similarity of the $N^{2}-N$ incorrect pairs(marked with superscript --). To be specific, the LLM is used to extract 10 types of discrete attributes of clothing (e.g., Category, FitType, CollarTechnique, IfTopTuckIn), which incorporates general in-context image knowledge from the language model [3, 5], eliminating semantic deficiencies and ambiguities present in vision. Then we assign a learnable embedding $f_{attr}$ to each discrete attribute. The image structure features $f_{img}$ are extracted by the twin ViT [19] encoder and concatenated with the attributes embeddings $f_{attr}$ to form the final latent structure embeddings $e$ after non-linear projection head. We optimize the InforNCE loss [15] over these latent structure embeddings $\left\{e_{itw},e_{std}\right\}$ from $\left\{x_{itw},x_{std}\right\}$ as:
$\small\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\text{log}\frac{\text{exp}(e_{itw,i}\odot e_{std}^{+}/\tau)}{\text{exp}(e_{itw,i}\odot e_{std}^{+}/\tau)+{\textstyle\sum_{e_{std}^{-}}}\text{exp}(e_{itw,i}\odot e_{std}^{-}/\tau)},$ (1)
where $\odot$ is cosine similarity between two vectors, $N$ is the batch size, and $\tau$ is a temperature scalar.
@@ 3.2 Retrieval-augmented SLLE
By integrating structural knowledge from StructureNet into the stable diffusion (SD), we have improved the style and structure of generated flat-lay clothing as shown in  Sec. 4.3. However, due to the limited spatial perception inherent in the SD [10, 37], it often struggles to accurately represent the length and contours of clothing items. This issue becomes particularly evident when the fit of garments on human bodies may not be well illustrated (Fig. 6), where we rely solely on soft embedding guidance $e_{std}$. Furthermore, since StructureNet $g_{\theta}$ is trained on a limited set of contrastive pairs, it may perform poorly with out-of-distribution samples, making it unreliable for real-world applications.
@@ 3.3 Omni-Level faithful garment generation
Structure faithfulness. Inspired by the global semantics control practice by IP-Adapter [69], we adopt a similar Embedding Prompt Adapter (EP-Adapter) to condition the high-level semantics of structure embeddings. While maintaining the text branch unchanged, fused structure embeddings $\hat{e}_{itw}$ after SLLE are fed into additional projection layers to generate key and value matrices for the structure representations. Two parallel cross-attention layers process the text modality and the structure embedding modality, with the attention results being summed to replace the original single text one: $\mathsf{Attention}(Q,K_{text},V_{text})+\mathsf{Attention}(Q,K_{emb},V_{emb})$. The EP-Adapter enhances the global/inner structural faithfulness of generation results by incorporating prior knowledge from LLM and structural training pairs as a soft constraint. Meanwhile, the contour landmark $\hat{L_{sil}}$ involves external spatial structural information. A Landmark Guider [28] with $4$ convolution layers ($4\times 4$ kernels, $2\times 2$ strides, $16,32,64,128$ channels) is incorporated to align the silhouette mask with noise $z_{t}$ as an explicit and hard constrain.
Pattern faithfulness. Inspired by success in human editing [28, 13], we introduce an additional U-Net encoder (i.e. ReferenceNet [28]) to precisely preserve the fine-grained details of clothing assets, which is isomorphic to the main generative U-Net (i.e. MainNet) and shares same initial parameter weights. The latent of in-the-wild garment image is passed through ReferenceNet to obtain the intermediate key and value features $\left\{z_{R}^{k},z_{R}^{v}\right\}\in\mathbb{R}^{N\times l\times d}$ in self-attention, which are concatenated with the features $\left\{z_{M}^{k},z_{M}^{v}\right\}\in\mathbb{R}^{N\times l\times d}$ from the MainNet along the $l$ dimension to obtain the final $\left\{z_{C}^{k},z_{C}^{v}\right\}\in\mathbb{R}^{N\times 2l\times d}$. Then we compute the self-attention on the concatenated features as:
@@
where $z_{M}^{q}$ represents Query features in self-attention from the MainNet. ReferenceNet plays an important role in preserving the texture of garments when it has complicated patterns or logo prints.
Decoding faithfulness. With ReferenceNet for pattern faithfulness, we still observe fine-grained distortion, because of the image decoding degradation caused by the high compression ratio of SDXL’s VAE, as shown in Fig. 6. Although the VAE in FLUX [1] greatly alleviates the loss by inflating latent channels, challenges intrinsic to the DiT framework [47], such as slow convergence, and high data requirements, hinder its widespread adoption in the community. To address this, we propose a versatile three-stage parameter gradual encoding adaption (PGEA) to align the SDXL UNet with the FLUX VAE. Specifically, we expand the channels in conv in and conv out layers (from 4 to 16) in U-Net to match FLUX VAE, enabling direct modification on the SD config file to load adapted weights. Stage I: we focus on distilling the knowledge from the original conv in layer to the adapted conv in layer. The input image is encoded through different VAEs and passed into the 4 and 16-channel conv in layers respectively, applying a reconstruction loss between their output to update the 16-channel conv in layer for fast adaption. Stage II: we train the UNet on a standard text-to-image task for 20,000 steps, where only the 16-channel conv in and conv out layers are updated. Stage III: we train and update the entire UNet on the standard text-to-image task for 200,000 steps. The complete training of PGEA lasts for 4 days on $8$ 80GB-A100 GPUs. It is noteworthy that the adapted general UNet with extremely low decoding loss can be applied to various downstream tasks.

### v2

@@ Abstract
Standard clothing asset generation involves restoring forward-facing flat-lay garment images displayed on a clear background by extracting clothing information from diverse real-world contexts, which presents significant challenges due to highly standardized structure sampling distributions and clothing semantic absence in complex scenarios. Existing models have limited spatial perception, often exhibiting structural hallucinations and texture distortion in this high-specification generative task. To address this issue, we propose a novel Retrieval-Augmented Generation (RAG) framework, termed RAGDiffusion, to enhance structure determinacy and mitigate hallucinations by assimilating knowledge from language models and external databases. RAGDiffusion consists of two processes: (1) Retrieval-based structure aggregation, which employs contrastive learning and a Structure Locally Linear Embedding (SLLE) to derive global structure and spatial landmarks, providing both soft and hard guidance to counteract structural ambiguities; and (2) Omni-level faithful garment generation, which introduces a coarse-to-fine texture alignment that ensures fidelity in pattern and detail components within the diffusing. Extensive experiments on challenging real-world datasets demonstrate that RAGDiffusion synthesizes structurally and texture-faithful clothing assets with significant performance improvements, representing a pioneering effort in high-specification faithful generation with RAG to confront intrinsic hallucinations and enhance fidelity.
@@ 1 Introduction
While the high quality of images generated by diffusion models [22, 47] has significantly lowered the barriers for individuals to engage in content creation, the generation of 2D standard clothing assets from in-the-wild data has been relatively underexplored [45]. Standard clothing asset generation refers to generating forward-facing flat-lay garment images [51] on a clear background by recovering clothing information from arbitrary real-world scenes (e.g., garments hung on hangers, worn by models, or casually laid on chairs in LABEL:{fig:_motivation}). Standard garments serve as a crucial intermediary variable, connecting various downstream applications such as e-commerce catalogs maintenance, garment design, product marketing and virtual fitting [21, 67, 66].
The garment generation task presents dual technical challenges distinct from conventional virtual try-on paradigms. First, standard clothing generation necessitates standard structure and frontal positioning while inherently limiting stylistic flexibility, constrained by the narrow manifold of high-specification garment distributions [64]. Second, the complexity of in-the-wild conditions involving occlusions, multi-layer outfits, truncations, or extreme lateral viewpoints notably exacerbates structural uncertainty and pattern ambiguity. Collectively, experimental evaluations reveal that existing methods [53, 23] fail to achieve practical usability thresholds with structural and texture distortion, particularly manifesting a pathological structure hallucinations in inaccurate structure, fitness and length estimation under challenging scenes (see  Fig. 3 and  Fig. 6), which we attribute to the limited spatial reasoning capacity inherent in Stable Diffusion (SD) architectures.
Despite its critical commercial relevance, the garment restoration problem remains insufficiently addressed in literature [59, 51, 53, 45]. Early GAN-based frameworks [59] suffer from pattern distortions constrained by their generative priors. Recent approaches [51, 53, 45] employing pre-trained SD exhibit several limitations: (i) inadequate handling of structural hallucinations in hard cases (occluded, multi-layer, etc.), (ii) commercially unfaithful logo/texture reproduction which is extremely sensitive to e-commerce sellers. In summary, no existing solution achieves the dual objectives of high-structural integrity and photorealistic texture synthesis under real-world complexity.
@@
We propose RAGDiffusion, as a Retrieval-Augmented Generation (RAG) [40] paradigm including two processes: information aggregation based on retrieval, and conditional generation with omni-level fidelity, to address the issue of structural and texture distortion respectively. (1) Retrieval-based Structure Aggregation: Our key insight lies in enhancing structural determinacy through the assimilating of external structural landmarks and world knowledge, thereby rectifying visual models’ misconceptions regarding garment structure, fitness and proportions. During RAG process, contrastive learning [20, 10] is firstly introduced to train a dual tower network to extract multi-modal structure embeddings from images of two branches as well as attributes derived from a frozen large language model (LLM) [2, 1]. Considering potential encoding bias in real-world data, we propose a structure retrieval named Structure Locally Linear Embedding (SLLE) to project the predicted structure embedding towards a standard manifold [43, 58] as well as offer a silhouette landmark. The remapped latent structure embedding and the landmarks provide soft and hard structure guidance respectively through Embedding Prompt Adapter and Landmark Guider to eliminate structural hallucinations. (2) Omni-level Faithful Garment Generation: RAGDiffusion establishes coarse-to-fine texture alignment for pattern faithfulness and detail faithfulness for generated garment. Specifically, we ensure the generated pattern matches the conditioning through ReferenceNet [23], a conditioning UNet isomorphic to the denoising UNet. However, problematic detail and logo artifacts persist partially due to resolution limitations of VAE decoders, rendering the generated images commercially unusable for e-commerce sellers, which is an under-addressed issue across prior research [51, 53, 45]. To mitigate reconstruction distortions inherent in original VAE [27], we propose Parameter Gradual Encoding Adaptation (PGEA) to adapt the SDXL [37] backbone to a more powerful VAE.
To the best of our knowledge, RAGDiffusion stands as a pioneering multi-modal RAG method to solve intrinsic hallucination and unfaithfulness during image synthesis. Furthermore, we discover two additional benefits brought by RAG: zero-shot generalizability for unseen scenarios via retrieval database expansion, and human-interpretable control mechanisms through landmark manipulation. The core contributions of this work are threefold:
@@
  • We propose RAGDiffusion, a RAG framework for clothing generation with representation learning and SLLE to offer structure determinacy and eliminate hallucination.
@@
  • We employ a coarse-to-fine texture alignment in RAGDiffusion, addressing pattern and detail aspects, ensuring faithfulness of clothing on this high-specification task.
@@
  • Comprehensive experiments on challenging in-the-wild sets demonstrate RAGDiffusion is capable of synthesizing both structure and texture faithful clothing assets, outperforming current methods by a substantial margin. Furthermore, the ablation study has validated the effectiveness of different parts of RAGDiffusion.
@@ 2 Related works
Retrieval-augmented generative models. Retrieval-augmented strategies leverage external databases to enhance the capabilities of generative models across a variety of tasks. For instance, the RETRO [5] modifies an LLM to effectively utilize external databases. In the realm of image synthesis, retrieval has been employed in both GANs [49, 7] and diffusion models [4, 46], accommodating 3D generation [44], video generation [62] and artistic styles [42]. The essence of these works lies in retrieving similar images to serve as mimetic references for generating specified content, particularly in cases where models are insufficiently trained. Our RAGDiffusion not only builds upon this generalizability through convenient retrieval expansion, but also aggregates structural information, thereby mitigating intrinsic hallucinations on the high-standard tasks.
@@
Controllable text-to-image diffusion models. To attain conditional control in text-to-image diffusion models, ControlNet [61], T2I-Adapter [33] and IP-Adapter [57] incorporate additional trainable modules to fuse conditions on feature maps. Additionally, recent investigations have utilized a variety of prompt engineering techniques [28, 56, 65] and implemented cross-attention constraints [9, 54, 25, 68] to facilitate more controllable generative processes. Furthermore, some studies further investigate multi-condition or multi-modal generation [24, 38, 69, 29]. However, these methods rely on the associative capabilities of stable diffusion, which cannot guarantee fidelity, particularly with the highly standardized sampling of standard garments.
Garment restoration. Standard clothing generation involves restoring flat-lay garment images on from real-world contexts. TileGAN [59] pioneered a two-stage GAN [18] framework, suffering from backbone capability limitations. Recent works including TryOffDiff [51], TryOffAnyone [53], and IGR [45] have adopted pretrained SD as base backbones, yielding obvious improvements through enhanced generative priors. Specifically, TryOffDiff utilizes SigLIP [60] to extract garment features; TryOffAnyone employs direct image concatenation for conditional generation; IGR implements ReferenceNet-inspired [23] architectures for pattern consistency. However, these methods exhibit three critical shortcomings: (1) They ignore structural hallucinations and distortions arising from real-world complexities (multi-layer, extreme viewpoints, etc.). (2) Their detail fidelity fails to meet commercial standards. (3) Adaptation to new garment categories/scenarios requires costly paired data and retraining. Our RAGDiffusion systematically resolves these limitations through RAG paradigm and a coarse-to-fine alignment pipeline.
@@ 3 Method
An overview of the RAGDiffusion is presented in Fig. 2. The backbone of RAGDiffusion employs the SDXL [41], with the preliminary detailed in Appendix. Given an in-the-wild clothing image $x_{itw}\in\mathbb{R}^{H\times W\times 3}$, RAGDiffusion is aimed to generate an authentic standard flat lay in-shop garment image $x_{std}$. A dual tower StructureNet extracts latent structure embeddings by contrastive learning as detailed in Sec. 3.1, and SLLE retrieves and fuses structure embeddings and landmarks as in Sec. 3.2. Coarse-to-fine alignment generation involves ReferenceNet,and PGEA for pattern, and decoding faithfulness described in  Sec. $3.$
@@ 3.1 Dual-tower embeddings extraction
Contrastive learning. Given a batch of NN (in-the-wild clothing $x_{itw}$, standard clothing $x_{std}$) pairs, StructureNet $g_{\theta}$ is trained to predict which of the $N\times N$ possible ($x_{itw}$, $x_{std}$) pairings across a batch actually occurred according to structure similarity. To do this, StructureNet learns a multi-modal embedding space by jointly training a dual tower encoder tailored for $x_{itw}$ and $x_{std}$ to maximize the cosine similarity of the embeddings of the NN real pairs (marked with superscript ++) in the batch while minimizing the similarity of the $N^{2}-N$ incorrect pairs(marked with superscript -). To be specific, the LLM is used to extract 10 types of discrete attributes of clothing (e.g., Category, FitType, CollarTechnique, IfTopTuckIn), which incorporates general in-context image knowledge from the language model [1, 2], eliminating semantic deficiencies and ambiguities present in vision. Then we assign a learnable embedding $f_{attr}$ to each discrete attribute. The image structure features $f_{img}$ are extracted by the twin ViT [14] encoder and concatenated with the attributes embeddings $f_{attr}$ to form the final latent structure embeddings $e$ after non-linear Resampler [57] head. We optimize the InforNCE loss [10] over these latent structure embeddings $\left\{e_{itw},e_{std}\right\}$ from $\left\{x_{itw},x_{std}\right\}$ as:
$\small\mathcal{L}=\frac{-1}{N}\sum_{i=1}^{N}\text{log}\frac{\text{exp}(e_{itw,i}\odot e_{std}^{+}/\tau)}{\text{exp}(e_{itw,i}\odot e_{std}^{+}/\tau)+{\textstyle\sum_{e_{std}^{-}}}\text{exp}(e_{itw,i}\odot e_{std}^{-}/\tau)},$ (1)
where $\odot$ is cosine similarity between two vectors, NN is the batch size, and $\tau$ is a temperature scalar.
@@ 3.2 Retrieval-augmented SLLE
By integrating structural knowledge from StructureNet into the SD, we have improved the style and structure of generated flat-lay clothing as shown in  Sec. 4.3. However, due to the limited spatial perception inherent in the SD [6, 30], it often struggles to accurately represent the length and contours of clothing, particularly in hard cases (occlusions, multi-layer, lateral viewpoints, etc.), as in Fig. 6. Furthermore, since StructureNet is trained on a limited set of contrastive pairs, it may perform poorly with out-of-distribution samples, making it unreliable for real-world applications.
@@
Structure conditioning. Inspired by IP-Adapter [57], we adopt a similar Embedding Prompt Adapter (EP-Adapter) to condition the high-level structure semantics. While maintaining the text branch unchanged, fused structure embeddings $\hat{e}_{itw}$ after SLLE are fed into additional projection layers to generate key and value matrices for the structure representations. Two parallel cross-attention layers process the text modality and the structure embedding modality, with results being summed to replace the original single text one: $\mathsf{Attention}(Q,K_{text},V_{text})+\mathsf{Attention}(Q,K_{emb},V_{emb})$. The EP-Adapter enhances the global/inner structural faithfulness by incorporating prior knowledge from LLM and structural training pairs as a soft constraint. Meanwhile, the contour landmark $\hat{L_{sil}}$ involves external spatial structural information. A Landmark Guider [23] with 44 convolution layers ($4\times 4$ kernels, $2\times 2$ strides, $16,32,64,128$ channels) is incorporated to align the silhouette mask with $z_{t}$ as an explicit and hard constrain.
@@ 3.3 Omni-Level faithful garment generation
Pattern faithfulness. Inspired by success in human editing [23, 8], we introduce an additional UNet encoder (i.e. ReferenceNet [23]) to precisely preserve the fine-grained details of clothing assets, which is isomorphic to the main generative UNet (i.e. MainNet) and shares same initial parameter weights. The latent of in-the-wild garment image is passed through ReferenceNet to obtain the intermediate key and value features $\left\{z_{R}^{k},z_{R}^{v}\right\}\in\mathbb{R}^{N\times l\times d}$ in self-attention, which are concatenated with the features $\left\{z_{M}^{k},z_{M}^{v}\right\}\in\mathbb{R}^{N\times l\times d}$ from the MainNet along the $l$ dimension to obtain the final $\left\{z_{C}^{k},z_{C}^{v}\right\}\in\mathbb{R}^{N\times 2l\times d}$. Then we compute the self-attention on the concatenated features as:
@@
where $z_{M}^{q}$ represents Query features in self-attention from MainNet. ReferenceNet plays an important role in texture preserving of garments when it has complicated patterns.
Detail faithfulness. The cloth restoration technique primarily serves e-commerce sellers for product catalog maintenance and marketing purposes. The requirement for brand logo fidelity is so stringent that existing methods’ results [51, 53, 45] are practically unusable even with ReferenceNet, partially because of the fine-grained logo decoding degradation caused by the high compression ratio of SDXL’s VAE, as shown in Fig. 6. Although the VAE in FLUX [17] greatly alleviates the loss by inflating latent channels, challenges intrinsic to the DiT framework [36], such as slow convergence, and high data requirements, hinder its widespread adoption in the community.
To address this, we propose a versatile three-stage parameter gradual encoding adaption (PGEA) to align the SDXL UNet with the FLUX VAE. Specifically, we expand the channels in conv in and conv out layers (from 4 to 16) in UNet to match FLUX VAE, enabling direct modification on the SD config file to load adapted weights. Stage I: we focus on distilling the knowledge from the original conv in layer to the adapted conv in layer. The input image is encoded through different VAEs and passed into the 4 and 16-channel conv in layers respectively, applying a reconstruction loss between their output to update the 16-channel conv in layer for fast adaption. Stage II: we train the UNet on a standard text-to-image task for 20,000 steps, where only the 16-channel conv in and conv out layers are updated. Stage III: we train and update the entire UNet on the standard text-to-image task for 200,000 steps. The complete training of PGEA lasts for 4 days on 88 H20 GPUs. It is noteworthy that the adapted general UNet with extremely low decoding loss can be applied to various downstream tasks. Crucially, despite marginal gains in quantitative metrics, RAGDiffusion’s breakthrough in brand logo fidelity makes it commercially viable, the first in industrial garment restoration.